A particle with a charge of -1.23 × 10-8 C is moving with instantaneous velocity v with rightwards arrow on top = (4.27 × 104 m/s)i with hat on top + (-3.79 × 104 m/s)j with hat on top. What is the magnitude of the force exerted on this particle by magnetic field B with rightwards arrow on top = (1.32 T)k with hat on top? (Give your answer in scientific notation using "N" as unit)
Answer #1

here,

charge on particle , Q = - 1.23 * 10^-8 C

the velocity , v = ( 4.27 i - 3.79 j) * 10^4 m/s

magnetic feild , B = 1.32 T k

the force exerted , F = q * ( v X B)

F = (-1.23 * 10^-8) * ( (( 4.27 i - 3.79 j) * 10^4) X 1.32 K)

F = ( 6.15 i + 6.93 j ) * 10^-4 N

the magnitude of force exerted , |F| = sqrt(6.15^2 + 6.93^2) * 10^-4 N

|F |= 9.27 * 10^-4 N